OpenStudy (anonymous):

Point A lies on the curve y = (2x +1)^3 + 1. If the y coordinate of A is 2, find the equations of the tangent and normal to the curve at A. If the tangent and normal at A meets the x-axis at B and C respectively, find the area of the triangle ABC.

OpenStudy (anonymous):

ok. step 1, get the value of x at point A. \[2=(2x+1)^3+1\]Solving for x:\[x=0\]Next, find the derivative:\[y'=6(2x+1)\]This will be the slope of your tangent line to the graph. Plugging in x=0 you get y'=6. The equation of a line through the point (0,2) with slope y'=6 is\[(y-2)=6(x-0)\]Or,\[y=6x+2\]

OpenStudy (anonymous):

The line normal to the curve at the point (0,2) will have a slope of \[m=\frac{-1}{6}\] the perpendicular line will have a slope that is the negative reciprocal of the tangent lines slope. Use this to get the equation of the normal line.

OpenStudy (anonymous):

Now what is point B? Well, if this is where the tangent line meets the x axis, it is the x-intercept of the line y=6x+2. So,it must be the point (-1/3, 0)

OpenStudy (anonymous):

You should e able to get the rest from here :)

OpenStudy (anonymous):

I got these answers too but was unsure as the answers are in fractions. My area of triangle ABC is 12 1/3 or 12.3333. Was it right?

OpenStudy (anonymous):

The normal line equation should be \[y=\frac{-1}{6}x+2\]So point C is at (12,0). Vector BC is <37/3,0> and vector BA is <-37/3,2>. Area of the triangle ABC is 37/3 or 12 1/3. Looks correct :)